OpenStudy (anonymous):

consider a swuare with side of lengh 's' and diagonal of length 'd'. write a formula expressing 'd' as a function of 's'...............is this correct: d(s) ?

OpenStudy (anonymous):

If diagonal d is going from on point corner of a square to the corner across from it, than we truly just have two triangles. Thinking back to Pythagorus, he stated that the hypotenuse of a right triangle (d) is related to its sides (s and s) threw his famous formula: \[a^2+b^2=c^2\] -- subbing in our values \[s^2+s^2=d^2\] -- which can be changed to \[d(s)=\sqrt{2s^2}\] -- or in another form \[d(s)= s*\sqrt{2}\]